### v1

# LLM2Vec: Large Language Models Are Secretly Powerful Text Encoders

Abstract
Large decoder-only language models (LLMs) are the state-of-the-art models on most of today’s NLP tasks and benchmarks. Yet, the community is only slowly adopting these models for text embedding tasks, which require rich contextualized representations. In this work, we introduce LLM2Vec, a simple unsupervised approach that can transform any decoder-only LLM into a strong text encoder. LLM2Vec consists of three simple steps: 1) enabling bidirectional attention, 2) masked next token prediction, and 3) unsupervised contrastive learning. We demonstrate the effectiveness of LLM2Vec by applying it to 3 popular LLMs ranging from 1.3B to 7B parameters and evaluate the transformed models on English word- and sequence-level tasks. We outperform encoder-only models by a large margin on word-level tasks and reach a new unsupervised state-of-the-art performance on the Massive Text Embeddings Benchmark (MTEB). Moreover, when combining LLM2Vec with supervised contrastive learning, we achieve state-of-the-art performance on MTEB among models that train only on publicly available data. Our strong empirical results and extensive analysis demonstrate that LLMs can be effectively transformed into universal text encoders in a parameter-efficient manner without the need for expensive adaptation or synthetic GPT-4 generated data.
1 Introduction
Text embedding models aim to encode the semantic content of natural language text in vector representations which then facilitate various natural language processing (NLP) tasks, such as semantic textual similarity, information retrieval, and clustering. For many years, the dominating paradigm for building such models relied on pre-trained bidirectional encoders or encoder-decoders such as BERT (Devlin et al., 2019) and T5 (Raffel et al., 2020), which are typically adapted for text embedding tasks by following a multi-step training pipeline consisting of weakly- and fully-supervised contrastive training (Ni et al., 2022; Li et al., 2023; Xiao et al., 2023, inter alia). Only recently, the community started to adopt decoder-only LLMs for embedding text (Muennighoff, 2022; Ma et al., 2023; Wang et al., 2023; Springer et al., 2024).
We speculate that the slow adoption of decoder-only LLMs for text embedding tasks is partly due to their causal attention mechanism, which inherently limits their ability to produce rich contextualized representations. At any given layer, causal attention limits token interactions, ensuring that the representation of a token at position $i$ is influenced solely by the representations of preceding tokens at positions $10,1,\dots,i-$ 1. Although this limitation is necessary for generative capabilities, it is sub-optimal for text embeddings as it prevents the representations from capturing information across the entire input sequence.
Figure 1: The 3 steps of LLM2Vec. First, we enable bidirectional attention to overcome the restrictions of causal attention (Bi). Second, we adapt the model to use bidirectional attention by masked next token prediction training (MNTP). Third, we apply unsupervised contrastive learning with mean pooling to learn better sequence representations (SimCSE).
Overcoming this architectural limitation of decoder-only LLMs for text embedding tasks is highly appealing as these models come with several advantages compared to their encoder-only counterparts. [^1] We acknowledge that there are also several challenges associated with the large size of these models and provide a discussion in Appendix A. During pre-training, decoder-only LLMs learn from all input tokens and not just a small percentage [^2] Encoder-only models are typically pre-trained by masking 15% of the tokens in the input sequence (Devlin et al., 2019)., which—given the same amount of training data—makes them much more sample-efficient than encoder-only models (Clark et al., 2020). Moreover, there exists a rich ecosystem around these models, with extensive tooling and well tested pre-training recipes, which has resulted in continuous improvement of these models by the community. Lastly, recent work on instruction fine-tuning and learning from human preferences has resulted in decoder-only LLMs that excel at instruction following (Wang et al., 2022b; Ouyang et al., 2022), making them an ideal choice for building universal text embedding models that generalize across a large variety of tasks using instructions.
In this work, we provide a simple unsupervised approach, termed LLM2Vec, which can be used to transform any pre-trained decoder-only LLM into a (universal) text encoder. As shown in Figure 1, LLM2Vec consists of three simple steps: 1) enabling bidirectional attention, 2) masked next token prediction, and 3) unsupervised contrastive learning. Crucially, LLM2Vec does not require any labeled data and is highly data- and parameter-efficient.
We apply LLM2vec to 3 decoder-only LLMs ranging from 1.3B to 7B parameters (S-LLaMA-1.3B, LLaMA-2-7B, and Mistral-7B) and evaluate the resulting models on word- and sequence-level tasks. On word-level tasks (chunking, named-entity recognition, and part-of-speech tagging), LLM2Vec-transformed models outperform strong encoder-only models by a large margin, demonstrating its effectiveness for producing rich contextualized token representations. On the Massive Text Embeddings Benchmark (MTEB), LLM2Vec-transformed models set a new state-of-the-art for unsupervised models, with our best model reaching a score of $56.8$. Additionally, we combine LLM2Vec with supervised contrastive training and achieve a new state-of-the-art performance among models that train only on publicly available data. Beyond our strong empirical results, we provide an extensive analysis of how LLM2Vec affects the representations of the underlying model and reveal an intriguing property of Mistral-7B, the currently best performing architecture on MTEB.
Overall, our work demonstrates that decoder-only LLMs are indeed capable of producing universal text embeddings and only very little adaptation is required to reveal this ability. Our code and pre-trained models are available at https://github.com/McGill-NLP/llm2vec.
2 LLM2Vec
2.1 Three simple ingredients
Enabling bidirectional attention (Bi)
The first step of the LLM2Vec approach is to replace the causal attention mask of decoder-only LLMs by an all-ones matrix (see Section B.1 for background on the self-attention mechanism). This gives each token access to every other token in the sequence, converting it into a bidirectional LLM (Devlin et al., 2019; Liu et al., 2019). However, it is not a priori clear why this should lead to better sequence representations. After all, the decoder-only LLM was not trained to attend to future tokens and therefore, this naive approach might even lead to worse representations. As we show, simply enabling bidirectional attention does indeed decrease embedding performance for most models. We can however easily adapt a model to make use of its bidirectional attention.
Masked next token prediction (MNTP)
We propose a simple strategy to make the model aware of its bidirectional attention by adapting it via masked next token prediction (MNTP). MNTP is a novel training objective that combines next token prediction with masked language modeling (Devlin et al., 2019). Given an arbitrary sequence $\mathbf{x}=(x_{1},x_{2},\ldots,x_{N})$ as input, we first mask a fraction of the input tokens and then train the model to predict the masked tokens based on the past and future context. Crucially, when predicting a masked token at position $i$, we compute the loss based on the logits obtained from the token representation at the previous position $i-1$, not the masked position itself (see Figure 1).
Unsupervised contrastive learning (SimCSE)
While the previous two steps of the LLM2Vec recipe can transform any decoder-only LLM into an encoder for word-level tasks, they might not be sufficient for sequence representations. Unlike bidirectional encoders that include a next sentence prediction objective in their pre-training objectives (Devlin et al., 2019), decoder-only LLMs are not explicitly trained to capture the context of the entire sequence. To fill this gap, we apply unsupervised contrastive learning via SimCSE (Gao et al., 2021). Specifically, given an input sequence, it is passed through the model twice with independently sampled dropout masks, resulting in two different representations for the same sequence. The model is trained to maximize the similarity between these two representations while minimizing the similarity with representations of other sequence in the batch. Crucially, this step does not require any paired data and can be applied using any collection of sequences. We use a pooling operation on the token representations to get the sequence representation (more details in Section 3.2).
The three steps combined – Bi + MNTP + SimCSE – constitute the LLM2Vec recipe.
2.2 Transforming decoder-only LLMs with LLM2Vec
Models
We experiment with 3 different decoder-only LLMs ranging from 1.3B to 7B parameters: Sheared-LLaMA-1.3B (S-LLaMA-1.3B, Xia et al., 2023), Llama-2-7B-chat (LLaMA-2-7B, Touvron et al., 2023), and Mistral-7B-Instruct-v0.2 (Mistral-7B, Jiang et al., 2023).
Training data
We perform both the MNTP and the unsupervised SimCSE step using data from English Wikipedia. We select data from Wikipedia as it was included in the pre-training mixture of all the models we experiment with. It is therefore fair to assume that these two adaptation steps are not teaching the model any new knowledge beyond how to attend to future tokens and how to construct sequence representations. Specifically, we use the Wikitext-103 dataset (Merity et al., 2017) for the MNTP step and a subset of Wikipedia sentences released by Gao et al. (2021) for the unsupervised SimCSE step.
Masked next token prediction
We follow established practice from the masked language modeling literature and randomly mask a fraction of the tokens from the input sequence (Devlin et al., 2019; Liu et al., 2019). We use the underscore (_) as the mask token, since the models we experiment with do not have a special token for masking. We fine-tune the model using LoRA (Hu et al., 2022) to predict the masked token using the representation of the previous token to maximally align our training objective with the pre-training setup of decoder-only LLMs. For all models, we trained for 1000 steps with a batch size of 32 on a single 80GB A100 GPU. For 7B models, this training takes only 90 minutes. We provide additional details of our training setup and hyperparameters in Section D.1.1.
Unsupervised contrastive learning
For the contrastive training, we apply the unsupervised SimCSE approach by Gao et al. (2021). The positive examples are constructed by applying the LLM’s dropout twice on the same input sequence, whereas other sequences in the batch act as in-batch negatives. We merge the MNTP LoRA weights into the base model and initialize new LoRA parameters before starting the SimCSE training, which ensures that the model retains the knowledge learned in the previous step. Similar to the MNTP step, we train for 1000 steps. For 7B models, this training takes 2.5 hours on a single 80GB A100 GPU with a batch size of 128. We provide additional details of our training setup and hyperparameters in Section D.1.2.
3 LLM2Vec-transformed models are strong unsupervised text embedders
3.1 Evaluation on word-level tasks
We start by evaluating on word-level tasks to demonstrate that LLM2Vec is successful at improving the contextual representations constructed by decoder-only LLMs.
(a) Chunking
(b) NER
(c) POS
Figure 2: Evaluation of LLM2Vec-transformed models on word-level tasks. Solid and dashed horizontal lines show the performance of Uni and DeBERTa-v3-large, respectively.
Setup
We evaluate on three word-level tasks: chunking, named-entity recognition (NER), and part-of-speech tagging (POS), using the CoNLL-2003 benchmark (Tjong Kim Sang & De Meulder, 2003). We embed each input sentence and train a task-specific linear classifier on top of the frozen representations. This is akin to the linear probing setup commonly used in the language model analysis literature (Belinkov, 2022). We compare the LLM2Vec-transformed models to DeBERTa-v3-large (He et al., 2023), the current state-of-the-art encoder-only model. Additional details about our setup are provided in Section D.1.3.
Results
Figure 2 shows the results of our evaluation (a detailed breakdown of the results is provided in Table 4). On each of the three tasks, constructing token representations with causal attention (Uni) already outperforms the encoder-only baseline. This is not surprising, given that the models we experiment with are significantly larger and have been pre-trained on more data. As expected, naively applying bidirectional attention dramatically hurts performance in most cases. Interestingly, for Mistral-7B, enabling bidirectional attention hurts performance much less compared to S-LLaMA-1.3B and LLaMA-2-7B. For NER, Mistral’s performance even improves by $0.6\%$ with bidirectional connections.
Next, we observe that for all models and tasks, further adapting via MNTP (i.e. LLM2Vec w/o SimCSE) improves performance. For instance, in the chunking task, we see improvements for S-LLaMA-1.3B (by $5\%$), LLaMA-2-7B (by $4\%$), and Mistral-7B (by $4\%$). Finally, combining MNTP with SimCSE, however, performs worse than just applying MNTP. This is expected for word-level tasks, as SimCSE adapts the representations for sequence-level tasks.
3.2 Evaluation on sequence-level tasks
(a) S-LLaMA-1.3B
(b) LLaMA-2-7B
(c) Mistral-7B
Figure 3: Unsupervised results on our 15-task subset of the MTEB dataset. We ablate three different pooling choices: EOS, mean pooling, and weighted mean pooling. LLM2Vec is compatible with all three approaches and works best with mean pooling.
Categories $\rightarrow$ Retr. Rerank. Clust. PairClass. Class. STS Summ. Avg
# of datasets $\rightarrow$ 15 4 11 3 12 10 1 56
Encoder-only models
BERT 10.59 43.44 30.12 56.33 61.66 54.36 29.82 38.33
BERT + SimCSE 20.29 46.47 29.04 70.33 62.50 74.33 31.15 45.45
S-LLaMA-1.3B
Uni + w. Mean 9.47 38.02 28.02 42.19 59.79 49.15 24.98 35.05
LLM2Vec (w/o SimCSE) 15.48 40.99 31.83 50.63 64.54 62.06 26.82 41.43
LLM2Vec 25.93 47.70 37.45 72.21 67.67 71.61 31.23 49.42
Echo embeddings 10.36 41.52 30.03 52.08 63.75 59.36 22.79 39.10
LLaMA-2-7B
Ui + w. Mean 15.16 46.94 36.85 61.41 69.05 63.42 26.64 44.54
LLM2Vec (w/o SimCSE) 19.86 44.74 35.31 61.60 67.94 66.74 26.83 45.70
LLM2Vec 36.75 52.95 40.83 77.89 71.57 76.41 31.38 55.36
Echo embeddings 16.16 46.84 34.25 63.54 69.82 67.95 25.57 45.36
Mistral-7B
Uni + w. Mean 10.43 45.11 35.82 60.28 71.14 58.59 26.57 42.46
LLM2Vec (w/o SimCSE) 19.74 50.43 40.06 70.95 72.51 71.90 27.84 49.43
LLM2Vec 38.05 53.99 40.63 80.94 74.07 78.50 30.19 56.80
Echo embeddings 22.68 51.07 36.78 75.87 72.69 73.60 29.54 50.26
Table 1: Unsupervised results on MTEB. We compare S-LLaMA-1.3B, LLaMA-2-7B, and Mistral-7B with and without LLM2Vec to the unsupervised BERT models of Gao et al. (2021) as well as Echo embeddings (Springer et al., 2024).
Next, we evaluate on the Massive Text Embedding Benchmark (MTEB), a collection of 7 diverse embedding task categories covering a total of 56 datasets (Muennighoff et al., 2023). To select the best-performing pooling method for each method, we perform ablations on a 15-task subset consisting of representative tasks from each of the MTEB categories. We provide additional details and justification for how we chose this subset in Section C.1.
Setup
Following previous work (Su et al., 2023; Wang et al., 2023; Springer et al., 2024), we evaluate with task-specific instructions. We use the same instructions as Wang et al. (2023) and only add instructions to queries. The instructions can be found in Table 10. For symmetric tasks, the same instruction will be used for the query and the document. When applying (weighted) mean pooling (Muennighoff, 2022), we exclude the instruction tokens.
As a baseline, we compare to the unsupervised BERT models obtained from Gao et al. (2021). Additionally, we compare to Echo embeddings, a concurrent approach by Springer et al. (2024), which we run with the same models and instructions (see Section E.1 for more details on our implementation of Echo embeddings). Echo duplicates the input and pools only over the second occurrence to address the limitation of causal attention.
Results on our 15-task subset of MTEB
Figure 3 shows the impact of various pooling methods for all three models on the subset of MTEB tasks. We can clearly observe that applying causal attention is sub-optimal when constructing text embeddings. The naive approach of applying the EOS pooling for models with causal attention is outperformed by (weighted) mean pooling. Enabling bidirectional attention without any training harms performance for S-LLaMA-1.3B and LLaMA-2-7B. Similar to our word-level results, the performance of Mistral-7B improves with bidirectional attention, even without any training.
For LLM2Vec-transformed models, applying MNTP training improves the performance of S-LLaMA-1.3B and Mistral-7B. Moreover, applying SimCSE further boosts the performance of S-LLaMA-1.3B, LLaMA-2-7B, and Mistral-7B by $49.8\%$, $23.2\%$, and $37.5\%$ compared to the best causal baseline on the MTEB subset. We conduct further ablations of each component of the LLM2Vec recipe in Section D.2.2 (Table 5).
Results on full MTEB
Table 1 shows the results of the best-performing models, which we select based on the ablation above, on the full MTEB dataset. After the first two steps of LLM2Vec—bidirectional attention and MNTP (LLM2Vec w/o SimCSE)—we observe a considerable improvement in performance for all three models (e.g., $16.4\%$ improvement for Mistral-7B).
When comparing to Echo embeddings, LLM2Vec (w/o SimCSE) [^3] We only directly compare the performance after the first two steps of LLM2Vec to Echo embeddings as applying SimCSE involves learning sequence representation, which makes the comparison unfair. leads to improved performance for S-LLaMA-1.3B and LLaMA-2-7B and performs almost on par for Mistral-7B. However, compared to Echo embeddings, LLM2Vec is much more efficient as Echo embeddings repeat the input and therefore double the sequence length which makes inference considerably slower (we provide a runtime comparison in Section E.2).
Adding the final step of the LLM2Vec recipe— unsupervised SimCSE—further boosts all three models by a large margin, making our LLM2Vec Mistral-7B state-of-the-art among all unsupervised models with a score of $56.80$.
Overall, our results highlight that LLM2Vec is successful at transforming decoder-only LLMs into strong text embedding models that outperform previous unsupervised approaches on the challenging MTEB leaderboard.
4 How does LLM2Vec affect a model?
4.1 LLM2Vec helps models to capture information from future tokens
To analyze the extent to which LLM2Vec-transformed models incorporate information from future tokens, we adopt the analysis of Springer et al. (2024) and test how good models are at judging the similarity of sentences that have different meanings but share the same prefix.
Setup
(a) S-LLaMA-1.3B
(b) Mistral-7B
Figure 4: Cosine similarity between query ($q$) and negative ($s^{-}$) as well as positive examples ($s^{+}$). Plots for LLaMA-2-7B and other approaches are shown in Appendix F.
We evaluate on a synthetic dataset collected by Springer et al. (2024), which consists of 35 sentence triples $\{(q_{i}$, $s^{+}_{i}$, $s^{-}_{i})\}_{i=1}^{35}$ with $q_{i}=(A_{i},B_{i})$, $s^{+}_{i}=(A_{i},C_{i})$, and $s^{-}_{i}=(A_{i},D_{i})$, where $B_{i}$ and $C_{i}$ have a similar meaning to each other but $B_{i}$ and $D_{i}$ don’t. We compute a sequence representation for each of these sentences by pooling only over the first part of the sentence, i.e., $A_{i}$. We then compute the cosine similarity between the resulting embeddings. A model that incorporates information from future tokens ($B_{i}$, $C_{i}$, or $D_{i}$) in the representations of the prefix $A_{i}$ should assign a higher similarity to the positive example.
Results
(a) S-LLaMA-1.3B
(b) LLaMA-2-7B
(c) Mistral-7B
Figure 5: Cosine similarities at different token positions at layers when comparing representations constructed with causal attention to those constructed with bidirectional attention (without training). Additional plots are shown in Appendix F.
Figure 4 shows the results of our analysis for S-LLaMA-1.3B and Mistral-7B. Results for LLaMA-2-7B, which show the same trends, and a comparison to Echo embeddings are provided in Appendix F. For S-LLaMA-1.3B, we observe that enabling bidirectional attention and training with the MNTP objective are sufficient to establish a clear separation between the positive and negative examples. For Mistral-7B, all setups lead to a larger cosine similarity between the query and positive than the query and negative examples.
4.2 Why does bidirectional attention without training work for Mistral models?
Our empirical results so far as well as the analysis above share an intriguing observation: enabling bidirectional attention works well for Mistral-7B, even without any training. Below, we investigate this surprising behavior by analyzing how bidirectional attention impacts the representations of a model.
Setup
We feed a single input sequence (a random paragraph from Wikipedia) to each model and compute the hidden representations of every token at every layer $l$ with causal ($\mathbf{H}^{c}_{l}$) and bidirectional attention ($\mathbf{H}^{bi}_{l}$). For every layer, we compute the cosine similarity between the representations constructed using causal and bidirectional attention, i.e., $\text{sim}(\mathbf{H}^{c}_{l},\mathbf{H}^{bi}_{l})$. For most layers, we expect this similarity to be low, as enabling bidirectional attention without any training should lead to substantially different representations.
Results
Figure 5 shows that as expected, for S-LLaMA-1.3B and LLaMA-2-7B, enabling bidirectional attention without training has a profound impact on the representations, leading to low cosine similarity across almost all layers and token positions. For Mistral-7B, on the other hand, the representations have very high cosine similarity throughout.
Based on these findings (we replicate these results for other inputs and other Mistral models in Appendix F) and the strong unsupervised results for Mistral-7B with bidirectional attention, we speculate that Mistral models are pre-trained with some form bidirectional attention, e.g., prefix language modeling (Raffel et al., 2020) – at least for some parts of its training. We leave a more detailed investigation of this intriguing behavior for future work.
5 Combining LLM2Vec with supervised contrastive learning
The final piece of our evaluation combines LLM2Vec with supervised contrastive learning.
5.1 LLM2Vec leads to strong performance on the MTEB leaderboard
Setup
For supervised training, we train on a replication of the public portion of the E5 dataset (Wang et al., 2023) curated by Springer et al. (2024). The dataset consists of approximately 1.5M samples and we provide details on its compilation in Section G.1. We follow standard practice and train the models with contrastive learning using hard negatives and in-batch negatives. We use LoRA fine-tuning for the supervised setting as well. The MNTP LoRA weights are merged into the base model, and the trainable LoRA weights are initialized with SimCSE weights. For LLM2Vec models that use just MNTP, the LoRA weights are randomly initialized. The training is performed for 1000 steps with a batch size of 512. We detail other hyperparameters in Section G.2.
For a fair comparison, we only compare to models trained on publicly available data and provide a comparison to the top entries on the MTEB leaderboard in Section G.3.
Results
Categories $\rightarrow$ Retr. Rerank. Clust. PairClass. Class. STS Summ. Avg
# of datasets $\rightarrow$ 15 4 11 3 12 10 1 56
Previous work w/ public data only
Instructor-xl 49.26 57.29 44.74 86.62 73.12 83.06 32.32 61.79
[…]
S-LLaMA-1.3B
Uni + w. Mean 51.02 54.65 39.90 83.57 71.64 82.16 30.05 60.44
LLM2Vec (w/o SimCSE) 51.44 55.38 43.57 86.20 72.21 83.58 30.01 61.85
LLM2Vec 51.49 55.58 43.24 85.80 72.98 83.62 30.12 61.96
LLaMA-2-7B
Uni + w. Mean 54.33 58.01 40.57 87.01 75.60 83.47 29.68 62.96
LLM2Vec (w/o SimCSE) 54.60 57.38 45.24 88.03 76.33 83.73 28.49 64.14
LLM2Vec 54.34 57.70 45.04 87.87 76.53 83.43 28.82 64.04
Mistral-7B
Uni + w. Mean 54.81 57.37 41.07 86.05 76.01 83.44 30.74 63.20
LLM2Vec (w/o SimCSE) 55.99 58.42 45.54 87.99 76.63 84.09 29.96 64.80
LLM2Vec 56.05 58.59 45.12 88.18 76.72 83.69 30.66 64.72
Table 2: Supervised results on full MTEB benchmark. The best performing LLM2Vec model achieves a new SOTA performance among models trained only on publicly available data.
Figure 6: Results on the 15-task subset of MTEB during training. For all three models, applying LLM2Vec before supervised training leads to better performance with fewer steps.
Table 2 shows the results of our evaluation. For all three models, transforming a model with LLM2Vec leads to improved performance over the strong Uni + weighted mean baseline. Performing unsupervised SimCSE is less crucial for supervised training, and even leads to slightly worse performance for LLaMA-2-7B and Mistral-7B compared to just performing the MNTP step of LLM2Vec. However, as we will show in Section 5.2, LLM2Vec with MNTP and SimCSE is much more sample-efficient, and therefore crucial in compute- or data-constrained settings. Notably, our best model, Mistral-7B (LLM2Vec w/o SimCSE) leads to a new state-of-the-art performance among models trained only on publicly available data.
5.2 LLM2Vec leads to more sample-efficient training
Setup
To demonstrate the sample-efficiency of LLM2Vec-transformed models, we save a checkpoint every 25 training steps and evaluate them on our 15-task subset of MTEB.
Results
As shown in Figure 6, LLM2Vec-transformed models achieve better performance earlier in training. This observation is consistent across all three models. For S-LLaMA-1.3B, the smallest of our three models, even performing just MNTP leads to considerably improved sample efficiency. These results are particularly encouraging for settings where it is hard to acquire high quality labeled data, a setting which we leave for future work.
6 Related Work
Supervised text encoders
Initially, supervised methods primarily relied on tasks such as natural language inference or sentence similarity to train BERT-like models for producing sentence embeddings (Conneau et al., 2017; Reimers & Gurevych, 2019). Subsequently, BERT-like models have also been adapted to tasks like retrieval (Karpukhin et al., 2020; Khattab & Zaharia, 2020). More recent methods have further improved these models through a complex multi-stage learning pipeline that consists of large-scale weakly supervised contrastive training followed by multi-task fine-tuning (Ni et al., 2022; Wang et al., 2022a; Li et al., 2023; Xiao et al., 2023) Recent approaches have focused on enhancing the generalization and transferability of text embeddings using instructions (Su et al., 2023; Asai et al., 2023).
Unsupervised text encoders
Another line of work has explored training text embedders in an unsupervised manner using only a set of unordered sentences. These unsupervised approaches typically create two different representations of the same sentence for contrastive learning. The methods vary in how they form these representations – perturbing the input sentence (Wu et al., 2020), or using different model instances (Carlsson et al., 2021). SimCSE (Gao et al., 2021), the approach used in this work, generates two representations of the same sentence by passing it through the model twice with different dropout masks.
Turning decoder-only LLMs into text encoders
While decoder-only LLMs have outperformed bidirectional encoders across a large variety of language understanding tasks (Brown et al., 2020; Touvron et al., 2023; Jiang et al., 2023, inter alia), their impact on sentence representation learning remains limited. The most common approaches in the literature use the final hidden state of the last token as the sentence embedding (Neelakantan et al., 2022; Ma et al., 2023; Wang et al., 2023). Very recent (and concurrent) works address the inherent limitation of LLMs for sentence representation – their causal attention mechanism.
Muennighoff et al. (2024) perform multi-task full fine-tuning using a combination of self-supervised language modeling with causal attention and supervised contrastive learning with bidirectional attention. In contrast, our proposed approach is much more computationally efficient, as it requires only parameter-efficient fine-tuning and 1000 gradient steps. Closest to our work is the concurrent work of Springer et al. (2024). They propose copying the input sequence and appending it to itself, which addresses the contextualization issue of causal attention as tokens in the copy of the input can now attend to “future” tokens in the previous sequence. While this performs well in practice, it significantly increases the computational cost at inference time, which can be particularly problematic for encoding longer documents. Our approach outperforms Springer et al. (2024), without inducing any additional computational overhead at inference time.
7 Conclusion
We present LLM2Vec, a strong unsupervised approach to transform any decoder-only LLMs into a (universal) text embedder. We perform an extensive evaluation on word- and sequence-level tasks and demonstrate the effectiveness of LLM2Vec in both unsupervised and supervised settings. Applying LLM2Vec to Mistral-7B achieves a new state-of-the-art performance on MTEB among unsupervised approaches as well as among supervised approaches that train only on publicly available data. Beyond our strong empirical contributions, we provide an extensive analysis of how LLM2Vec impacts the underlying model and reveal an intriguing property of Mistral-7B, which explains its strong out of-the-box performance with bidirectional attention. The simplicity of our approach, as well as its compute and sample efficiency, makes LLM2vec a promising solution for low-resource and compute-constrained scenarios and opens up several interesting avenues for future work.
Acknowledgements
We thank the members of SR’s research group for providing feedback throughout the project. Furthermore, we thank Jacob Mitchell Springer for providing the supervised training data used in Springer et al. (2024). PB is supported by the Mila-Intel Grant program. MM is partly funded by the Mila P2v5 Technology Maturation Grant and the Mila-Samsung grant. SR is supported by a Facebook CIFAR AI Chair and NSERC Discovery Grant program.
Appendix A Limitations
Large size of decoder-only LLMs
Recent years have seen a increasing trend towards training very large decoder-only LLMs, with model sizes up to 540B parameters (Brown et al., 2020; Chowdhery et al., 2023). The parameter size of the model has a direct impact on the training and inference latency. Additionally, the large output embedding dimension of these models (e.g., 4096 for Mistral-7B compared to 768 for BERT) also makes them more memory and compute intensive for creating vector indexes for large document collections. While some of these limitations can be offset by recent advances in improving the training and inference efficiency of large models (Hu et al., 2022; Dao, 2024), these techniques can technically be applied to smaller bidirectional models as well.
The advantages of small bidirectional encoders come at the cost of complex and computationally intensive training regimes (Li et al., 2023; Xiao et al., 2023; Li et al., 2023). In contrast, decoder-only models are much more sample-efficient and do not require large-scale contrastive pre-training (Wang et al., 2023). Moreover, the instruction following capabilities of decoder-only models make them strong contenders for building text embedding models that generalize to a wide range of tasks and domains without the need for expensive adaptation.
While smaller models can be more practical for some applications, the sample-efficiency, the instruction following capabilities, and the widespread use of these models in the community motivates the need to explore the potential of decoder-only LLMs for text embedding tasks.
Data contamination from pre-training
As our supervised data contains train splits of publicly available datasets, there is an extremely low chance of test set contamination with the MTEB benchmark. However there is a possibility of contamination from the pre-training data of LLaMA-2-7B and Mistral-7B models (S-LLaMA-1.3B was distilled from LLaMA-2-7B). As the complete details of the pre-training data are not publicly available, we cannot be certain about the extent of contamination. However, to reliably compare with other works, we stick to our choice of model and evaluation benchmark. We leave it to future work to investigate the performance of these models on newly designed benchmarks that are not part of their pre-training data.
Extending to other languages
In this work, we have implemented and evaluated our proposed methodology – LLM2Vec – using only English text corpora and benchmarks. However, the methodology is language-agnostic and can be easily extended to other languages using just unstructured text collections. We leave it to future work to investigate the performance of LLM2Vec on other languages.
Appendix B Background
B.1 Self-attention
The self-attention mechanism is a crucial component of decoder-only LLMs Cheng et al. (2016); Lin et al. (2017); Vaswani et al. (2017); Paulus et al. (2018). Given a sequence of $N$ tokens, the token representation at any given transformer layer $(\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{N})$ with $\mathbf{x}_{i}\in\mathbf{R}^{d}$ are stacked into a matrix $\mathbf{X}\in\mathcal{R}^{N\times d}$. Given this matrix, the self-attention mechanism computes the query, key, and value matrices $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathcal{R}^{N\times p}$ via a learned linear transformation.
$\mathbf{Q}$ $=X\mathbf{W}^{Q},$ (1)
$\mathbf{K}$ $=X\mathbf{W}^{K},$ (2)
$\mathbf{V}$ $=X\mathbf{W}^{V}~{}.$ (3)
The output of the self-attention layer is then computed as a linear combination of the values, weighted by the normalized inner product between keys and queries:
$\mathbf{O}=\text{softmax}\left(\frac{\mathcal{M}_{\{j\leq i\}}\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d}}\right)\mathbf{V}~{}.$ (4)
This output is then passed through a feed-forward network and added to the residual stream to obtain the token representations at the next layer. Crucially, in the case of decoder-only LLMs, the attention mask $\mathcal{M}_{\{j\leq i\}}$ prevents accessing token embeddings to the right of the current token.
B.2 Contrastive learning
Contrastive learning is a popular paradigm to learn text representations Karpukhin et al. (2020); Gao et al. (2021); Su et al. (2023); Wang et al. (2023); Springer et al. (2024). In the supervised setup, we have a set of positive pairs $\mathcal{D}=\{(q_{i},d^{+}_{i})\}_{i=1}^{n}$, and a set of negative documents that can include hard or in-batch negatives. The model is trained to maximize the similarity (i.e., usually cosine similarity) of positive pairs and minimize the similarity of negative pairs, i.e., we optimize the following objective:
$\mathcal{L}=\frac{e^{\lambda s(q,d^{+})}}{e^{\lambda s(q,d^{+})}+\sum_{d^{-}\in N}{e^{\lambda s(q,d^{-})}}}~{},$ (5)
where $s$ is a similarity metric, $\lambda$ a temperature value, and $N$ all the negative documents for query $q$.
Unsupervised contrastive learning
In unsupervised contrastive learning, no positive or hard negative pairs are available. Most unsupervised approaches construct two different representation for the same sample, using either model or input perturbations. SimCSE (Gao et al., 2021), the unsupervised approach used in this work, creates two different representations of the same input by using independently sampled dropout masks in the intermediate model representations and train the model with in-batch negatives.
Appendix C Massive Text Embeddings Benchmark (MTEB)
C.1 MTEB subset details
MTEB consists of diverse small and large embedding tasks. To speed up the evaluation [^4] Full evaluation on MTEB takes more than 40h for Mistral-7B on 8x A100 GPUs., we consider a representative subset of 15 tasks from MTEB for our analyses, presented in Table 3. To make sure that our ablation and analyses are not biased towards one specific category or task, this subset includes tasks from each category with almost the same proportion compared to the full MTEB [^5] Since the MTEB’s SummEval category includes only one dataset, we skip this category in our small-scale evaluation..
Category Dataset
Retrieval (3) SciFact
ArguAna
NFCorpus
Reranking (2) StackOverflowDupQuestions
SciDocsRR
Clustering (3) BiorxivClusteringS2S
MedrxivClusteringS2S
TwentyNewsgroupsClustering
Pair Classification (1) SprintDuplicateQuestions
Classification (3) Banking77Classification
EmotionClassification
MassiveIntentClassification
STS (3) STS17
SICK-R
STSBenchmark
SummEval (0) -
Overall 15 datasets
Table 3: Subset of MTEB tasks used for our ablations and analysis.
C.2 MTEB instructions
When evaluating on MTEB, we use the same instructions as Wang et al. (2023). The list of instructions for each task is listed in Table 10.
Appendix D Details on unsupervised results
D.1 Training details
D.1.1 MNTP training details
The second step of LLM2Vec includes MNTP training. We follow established practice from the encoder-only literature for choosing our masking strategy. For example, (Devlin et al., 2019) mask $15\%$ of the tokens in the input. $10\%$ of the masked tokens are then replaced with a random token from the vocabulary, while another $10\%$ are unmasked again, but still considered when computing the loss. As another example, RoBERTa (Liu et al., 2019) also masks $15\%$ of the input tokens but applies no further post-processing to the masked tokens.
For our models, we perform a hyperparameter search to select the percentage of the masked tokens in a sequence choosing from $20\%,40\%,60\%,80\%$, and $90\%$. For each model, we take the best setup (i.e., masking probability and BERT vs. RoBERTa approach) based on the performance on SICK-R (Agirre et al., 2014) task from the MTEB dataset. This results in the following choices: for S-LLaMA-1.3B and LLaMA-2-7B, we apply BERT’s masking strategy with masking probability of $20\%$. For Mistral-7B, we apply RoBERTa’s masking strategy with probability of $80\%$.
We train all the models for 1000 steps with LoRA $r=16$ and $\alpha=32$, and we follow the same training parameters as RoBERTa MNTP training. When training large 7B models, we apply brain floating point (bfloat16) quantization, as well as FlashAttention-2 (Dao, 2024) and gradient checkpointing.
D.1.2 SimCSE training details
The last step of LLM2Vec involves unsupervised contrastive learning with SimCSE. Our initial experiments indicated that the low value of dropout probability ($0.1$) typically used by bidirectional encoders (Gao et al., 2021) does not lead to optimal performance for larger decoder-only LLMs. Therefore, we use a higher dropout probability of $0.3$ for all models.
Similar to MNTP, we train all models with LoRA $r=16$ and $\alpha=32$ for 1000 steps. For LLaMA-2-7B and Mistral-7B, we train with a batch size of 128. For S-LLaMA-1.3B, we use a batch size of 32. Additionally, when training LLaMA-2-7B and Mistral-7B, we apply brain floating point (bfloat16) quantization, FlashAttention-2 (Dao, 2024), and gradient checkpointing.
D.1.3 Word-level training details
We evaluate on three popular word embedding tasks: chunking, named-entity recognition (NER), and part-of-speech (POS) tagging. We train a linear classifier using dropout with a dropout probability of 0.1 on top of the frozen representations obtained from the last hidden layer of a model.
We use data from CoNLL-2003, consisting of roughly 14,000 training, 3,250 validation, and 3,450 test samples (Tjong Kim Sang & De Meulder, 2003). We train the classifier for 1,500 steps with a learning rate of $0.0005$ and batch size 8.
As the models we experiment with have sub-token based vocabularies, we calculate the embedding of a word by averaging the representations of all its sub-tokens. For example, for a sentence “$w_{1}\;w_{2}\;w_{3}$” which is tokenized as “$\text{BOS}\;\;t_{11}t_{12}\;\;t_{21}t_{22}t_{23}\;\;t_{31}$”, the representation of $w_{1}$, $w_{2}$, and $w_{3}$ will be computed as
$e_{1}=\frac{1}{2}\left(e_{11}+e_{12}\right),\;\;\;e_{2}=\frac{1}{3}\left(e_{21}+e_{22}+e_{23}\right),\;\;\;e_{3}=e_{31}.$
Here, $e_{.}$ is the final representation of token $t_{.}$ or word $w_{.}$. Moreover, for the models that have gone through MNTP, we calculate the representation based on sub-tokens of the previous word. Using the same example as above, for models trained with MNTP, the representation of words $w_{1}$, $w_{2}$, and $w_{3}$ will be computed as:
$e_{1}=\frac{1}{2}\left(e_{\text{BOS}}+e_{11}\right),\;\;\;e_{2}=\frac{1}{3}\left(e_{12}+e_{21}+e_{22}\right),\;\;\;e_{3}=e_{23}.$
D.2 Additional results
D.2.1 Word-level task results
We present the detailed breakdown of the performance of LLM2Vec-transformed models on the word-level tasks in Table 4. Our results show that applying MNTP training to decoder-only LLMs helps them take advantage of the enabled bidirectional attention which boosts their performance on word-level tasks.
Model Chunking NER POS tagging
Encoder-only
BERT-large 71.77 90.09 75.12
DeBERTa-large 85.74 94.97 86.49
S-LLaMA-1.3B
Uni 86.10 96.09 90.89
Bi 76.50 92.17 89.18
Bi + MNTP 90.51 96.59 92.04
Bi + SimCSE 75.93 91.45 89.22
Bi + MNTP + SimCSE (LLM2Vec) 89.33 95.90 90.38
LLaMA-2-7B
Uni 88.23 96.59 91.53
Bi 78.24 92.31 90.62
Bi + MNTP 91.61 97.16 92.61
Bi + SimCSE 77.75 91.96 90.48
Bi + MNTP + SimCSE (LLM2Vec) 89.66 96.05 90.53
Mistral-7B
Uni 87.53 96.52 90.86
Bi 85.66 97.14 90.70
Bi + MNTP 91.17 97.18 92.35
Bi + SimCSE 86.91 97.15 92.12
Bi + MNTP + SimCSE (LLM2Vec) 90.69 96.87 92.08
Table 4: Unsupervised results on the word-level tasks for different models.
D.2.2 Sentence-level task results
Table 5 presents the results on MTEB subset for all models across different pooling methods. Results show that While weighted mean works the best for causal (i.e., Uni) models, mean pooling performs the best for LLM2Vec approach.
Model EOS Mean W. mean
S-LLaMA-1.3B
Uni 27.72 33.03 34.99
Bi 21.16 30.26 30.20
Bi + MNTP 29.16 42.10 38.67
Uni + SimCSE 37.44 44.95 47.13
Bi + SimCSE 40.43 44.46 44.83
Bi + MNTP + SimCSE (LLM2Vec) 45.57 52.40 50.23
LLaMA-2-7B
Uni 33.23 45.83 47.85
Bi 34.47 38.22 37.50
Bi + MNTP 32.66 48.00 44.30
Uni + SimCSE 38.47 52.03 53.55
Bi + SimCSE 40.37 44.13 44.08
Bi + MNTP + SimCSE (LLM2Vec) 50.61 58.97 55.75
Mistral-7B
Uni 22.12 43.00 44.01
Bi 25.17 50.07 45.20
Bi + MNTP 26.54 53.89 48.93
Uni + SimCSE 34.60 52.04 53.95
Bi + SimCSE 49.73 60.29 56.56
Bi + MNTP + SimCSE (LLM2Vec) 53.67 60.50 57.55
Table 5: Unsupervised results on MTEB subset for different models.
In Table 11, we additionally present a breakdown of the unsupervised performance of LLM2Vec-transformed models on MTEB.
Appendix E Comparison with Echo embedding
E.1 Reproducibility
Concurrent to our work, Springer et al. (2024) proposed Echo embeddings, a simple approach to convert decoder-only LLMs into text embedders by copying the input sequence and appending it to itself. For evaluation, they follow a prompt sampling procedure for the task instruction. However, they report that the exact wording or template used as a prompting strategy does not have a strong effect on the performance.
For a fair comparison to our proposed models, we implement Echo embeddings using the instructions in our evaluation setup (Section C.2). To do a sanity check on our implementation, as well as to see the impact of exact wording of instructions, we evaluate Echo embedding on the same subset of 26 MTEB tasks that was chosen in their work. We run this evaluation using the Mistral-7B-Instruct-v0.1 model to ensure that the results are directly comparable to theirs.
The unsupervised Echo model based on our implementation and instructions achieved a score of $55.22$ on the 26 task subset, whereas their reported score is $55.07$. This result validates our implementation and confirms an observation made by Springer et al. (2024) – the exact wording or template used does not have a strong effect on the performance.
E.2 Efficiency
In Table 6, we report the approximate evaluation time it took (in hours) to evaluate each of the models on MTEB using 8x 80GB A100 GPUs. Given that Echo embeddings rely on copying the input text, evaluation takes much longer compared to our approach. We note that the increased inference time of Echo embeddings is especially problematic for the encoding of large retrieval corpora in MTEB benchmark.
Model LLM2Vec Echo embeddings
S-LLaMA-1.3B $\approx 30$ hrs $\approx 40$ hrs
LLaMA-2-7B $\approx 42$ hrs $\approx 63$ hrs
Mistral-7B $\approx 44$ hrs $\approx 64$ hrs
Table 6: Evaluation time of Echo Embeddings compared to LLM2Vec in hours on 8x 80GB A100 GPUs.
Appendix F More analysis results
Data used for our analysis
Table 7 shows examples of the data used for our cosine similarity analysis in Section 4.
$q$: the query
$s^{+}$: the positive sample
$s^{-}$: the negative sample
She loves to travel in summer, especially to cold destinations, avoiding hot and crowded places.
She loves to travel in summer, specifically to chilly locations, steering clear of warm, populous areas.
She loves to travel in summer, but prefers to visit hot and bustling tourist spots.
The cat often sits by the window, dreaming of chasing birds and enjoying the warm sunshine.
The cat often sits by the window, imagining bird pursuits and basking in the sunlight.
The cat often sits by the window, but is too lazy to dream of chasing anything.
He reads books every night, finding solace in fiction and escaping from the stresses of daily life.
He reads books every night, seeking comfort in stories and evading everyday tensions.
He reads books every night, yet he feels that non-fiction is more engaging and informative.
She paints landscapes on weekends, expressing her creativity through vibrant colors and abstract forms.
She paints landscapes on weekends, showcasing her artistic flair with lively hues and unconventional shapes.
She paints landscapes on weekends, preferring realistic and detailed depictions of nature.
Table 7: Toy data used for the analysis in Section 4. These sentences were originally collected by Springer et al. (2024).
Cosine similarity analysis
Figure 7 provide cosine similarity results for all three models. In addition to our LLM2Vec-transformed models, we also provide results for Echo embeddings.
F.1 Additional plots
(a) S-LLaMA-1.3B
(b) LLaMA-2-7B
(c) Mistral-7B
Figure 7: Cosine similarity between query and negative as well as positive examples for S-LLaMA-1.3B, LLaMA-2-7B, and Mistral-7B.
(a) S-LLaMA-1.3B
(b) LLaMA-2-7B
(c) Mistral-7B
(d) S-LLaMA-1.3B
(e) LLaMA-2-7B
(f) Mistral-7B
Figure 8: Cosine similarities at different token positions at layers when comparing representations constructed with causal attention to those constructed with bidirectional attention (without training).
(a) Mistral-7B-v0.1
(b) Mistral-7B-Instruct-v0.1
(c) Mistral-7B-Instruct-v0.2
Figure 9: Cosine similarities at different token positions at layers when comparing representations of Mistral models constructed with causal attention to those constructed with bidirectional attention (without training).
Representation analysis
Figure 8 provides additional plots for the representation analysis using two different Wikipedia paragraphs. The trends closely follow those reported in Section 4. Figure 9 shows that the same behavior we observe for Mistral-7B-instruct-v0.2 also holds true for other variants of the Mistral-7B model. We take this as additional evidence that the Mistral-7B base model was trained with some form of bidirectional attention.
Appendix G Details on supervised results
G.1 E5 dataset
The dataset consists of ELI5 (sample ratio 0.1) (Fan et al., 2019), HotpotQA (Yang et al., 2018), FEVER (Thorne et al., 2018), MIRACL (Zhang et al., 2023), MS-MARCO passage ranking (sample ratio 0.5) and document ranking (sample ratio 0.2) , NQ (Karpukhin et al., 2020), NLI (Gao et al., 2021), SQuAD (Rajpurkar et al., 2016), TriviaQA (Joshi et al., 2017), Quora Duplicate Questions (sample ratio 0.1) (DataCanary et al., 2017), Mr- TyDi (Zhang et al., 2021), DuReader (He et al., 2018), and T2Ranking (sample ratio 0.5) (Xie et al., 2023). The instruction used for each dataset can be found in Table 8.
Dataset Instruction(s)
NLI
Given a premise, retrieve a hypothesis that is entailed by the premise
Retrieve semantically similar text
DuReader
Given a Chinese search query, retrieve web passages that answer the question
ELI5
Provided a user question, retrieve the highest voted answers on Reddit ELI5 forum
FEVER
Given a claim, retrieve documents that support or refute the claim
HotpotQA
Given a multi-hop question, retrieve documents that can help answer the question
MIRACL
Given a question, retrieve Wikipedia passages that answer the question
MrTyDi
Given a question, retrieve Wikipedia passages that answer the question
MSMARCO Passage
Given a web search query, retrieve relevant passages that answer the query
MSMARCO Document
Given a web search query, retrieve relevant documents that answer the query
NQ
Given a question, retrieve Wikipedia passages that answer the question
QuoraDuplicates
Given a question, retrieve questions that are semantically equivalent to the given question
Find questions that have the same meaning as the input question
SQuAD
Retrieve Wikipedia passages that answer the question
T2Ranking
Given a Chinese search query, retrieve web passages that answer the question
TriviaQA
Retrieve Wikipedia passages that answer the question
Table 8: Instructions used for each of the E5 datasets.
G.2 Training details
All models are trained with LoRA $r=16$ and $\alpha=32$, brain floating point (bfloat16) quantization, gradient checkpointing, and FlashAttention-2 (Dao, 2024) to optimize GPU memory consumption. We train on 8 NVIDIA A100 GPUs with an effective batch size of 512 for 1000 steps using a maximum sequence length of 512 tokens. We use the Adam optimizer with a learning rate of $0.0002$ and a linear learning rate warm-up for the first 300 steps.
G.3 Results
[…]
Rank Model Size Public Embed. Retr. Rerank. Clust. PairClass. Class. STS Summ. Avg
(GB) Data Dim. 15 4 11 3 12 10 1 56
1 SFR-Embedding-Mistral 14.22 $\times$ 4096 59.00 60.64 51.67 88.54 78.33 85.05 31.16 67.56
2 voyage-lite-02-instruct 2.45 $\times$ 1024 56.60 58.24 52.42 86.87 79.25 85.79 31.01 67.13
3 GritLM-7B 14.48 $\times$ 4096 57.41 60.49 50.61 87.16 79.46 83.35 30.37 66.76
4 e5-mistral-7b-instruct 14.22 $\times$ 4096 56.89 60.21 50.26 88.34 78.47 84.63 31.40 66.63
5 GritLM-8x7B 93.41 $\times$ 4096 55.09 59.80 50.14 84.97 78.53 83.26 29.82 65.66
6 LLM2Vec (w/o SimCSE) 14.22 $\checkmark$ 4096 55.99 58.42 45.54 87.99 76.63 84.09 29.96 64.80
6 LLM2Vec 14.22 $\checkmark$ 4096 56.05 58.59 45.12 88.18 76.72 83.69 30.66 64.72
7 echo-mistral-7b-instruct-lasttoken 14.22 $\checkmark$ 4096 55.52 58.14 46.32 87.34 77.43 82.56 30.73 64.68
8 mxbai-embed-large-v1 0.67 $\times$ 1024 54.39 60.11 46.71 87.20 75.64 85.00 32.71 64.68
9 UAE-Large-V1 1.34 $\times$ 1024 54.66 59.88 46.73 87.25 75.58 84.54 32.03 64.64
10 text-embedding-3-large - $\times$ 3072 55.44 59.16 49.01 85.72 75.45 81.73 29.92 64.59
Table 9: Top-10 models on the MTEB leaderboard as of 2024-03-29. LLM2Vec achieves the 6th rank overall, and the top rank among models trained with only publicly available data.
Task Name Instruction
AmazonCounterfactualClassif.
Classify a given Amazon customer review text as either counterfactual or not-counterfactual
AmazonPolarityClassification
Classify Amazon reviews into positive or negative sentiment
AmazonReviewsClassification
Classify the given Amazon review into its appropriate rating category
Banking77Classification
Given a online banking query, find the corresponding intents
EmotionClassification
Classify the emotion expressed in the given Twitter message into one of the six emotions: anger, fear, joy, love, sadness, and surprise
ImdbClassification
Classify the sentiment expressed in the given movie review text from the IMDB dataset
MassiveIntentClassification
Given a user utterance as query, find the user intents
MassiveScenarioClassification
Given a user utterance as query, find the user scenarios
MTOPDomainClassification
Classify the intent domain of the given utterance in task-oriented conversation
MTOPIntentClassification
Classify the intent of the given utterance in task-oriented conversation
ToxicConversationsClassif.
Classify the given comments as either toxic or not toxic
TweetSentimentClassification
Classify the sentiment of a given tweet as either positive, negative, or neutral
ArxivClusteringP2P
Identify the main and secondary category of Arxiv papers based on the titles and abstracts
ArxivClusteringS2S
Identify the main and secondary category of Arxiv papers based on the titles
BiorxivClusteringP2P
Identify the main category of Biorxiv papers based on the titles and abstracts
BiorxivClusteringS2S
Identify the main category of Biorxiv papers based on the titles
MedrxivClusteringP2P
Identify the main category of Medrxiv papers based on the titles and abstracts
MedrxivClusteringS2S
Identify the main category of Medrxiv papers based on the titles
RedditClustering
Identify the topic or theme of Reddit posts based on the titles
RedditClusteringP2P
Identify the topic or theme of Reddit posts based on the titles and posts
StackExchangeClustering
Identify the topic or theme of StackExchange posts based on the titles
StackExchangeClusteringP2P
Identify the topic or theme of StackExchange posts based on the given paragraphs
TwentyNewsgroupsClustering
Identify the topic or theme of the given news articles
SprintDuplicateQuestions
Retrieve duplicate questions from Sprint forum
TwitterSemEval2015
Retrieve tweets that are semantically similar to the given tweet
TwitterURLCorpus
Retrieve tweets that are semantically similar to the given tweet
AskUbuntuDupQuestions
Retrieve duplicate questions from AskUbuntu forum
MindSmallReranking
Retrieve relevant news articles based on user browsing history
SciDocsRR
Given a title of a scientific paper, retrieve the titles of other relevant papers
StackOverflowDupQuestions
Retrieve duplicate questions from StackOverflow forum
ArguAna
Given a claim, find documents that refute the claim
ClimateFEVER
Given a claim about climate change, retrieve documents that support or refute the claim
CQADupstackRetrieval
Given a question, retrieve detailed question descriptions from Stackexchange that are duplicates to the given question
DBPedia
Given a query, retrieve relevant entity descriptions from DBPedia
FEVER
Given a claim, retrieve documents that support or refute the claim
FiQA2018
Given a financial question, retrieve user replies that best answer the question
HotpotQA
Given a multi-hop question, retrieve documents that can help answer the question
MSMARCO
Given a web search query, retrieve relevant passages that answer the query
NFCorpus
Given a question, retrieve relevant documents that best answer the question
NQ
Given a question, retrieve Wikipedia passages that answer the question
QuoraRetrieval
Given a question, retrieve questions that are semantically equivalent to the given question
SCIDOCS
Given a scientific paper title, retrieve paper abstracts that are cited by the given paper
SciFact
Given a scientific claim, retrieve documents that support or refute the claim
Touche2020
Given a question, retrieve detailed and persuasive arguments that answer the question
TRECCOVID
Given a query on COVID-19, retrieve documents that answer the query
STS*
Retrieve semantically similar text.
BUCC/Tatoeba
Retrieve parallel sentences.
SummEval
Given a news summary, retrieve other semantically similar summaries
Table 10: Instructions used for evaluation on the MTEB benchmark. “STS*” refers to all the STS tasks.
Task S-LLaMA-1.3B LLaMA-2-7B Mistral-7B
AmazonCounterfactualClassification 72.93 76.91 76.94
AmazonPolarityClassification 74.28 79.05 85.29
AmazonReviewsClassification 36.14 40.08 47.09
ArguAna 43.64 47.09 51.00
ArxivClusteringP2P 42.91 47.81 47.56
ArxivClusteringS2S 35.20 40.53 39.92
AskUbuntuDupQuestions 52.70 55.56 58.60
BIOSSES 75.12 82.41 83.29
Banking77Classification 79.00 84.65 86.16
BiorxivClusteringP2P 35.02 38.12 36.14
BiorxivClusteringS2S 27.21 31.25 30.26
CQADupstackRetrieval 18.50 30.78 33.37
ClimateFEVER 18.95 20.67 22.97
DBPedia 13.21 25.81 25.48
EmotionClassification 42.85 46.58 48.88
FEVER 16.96 43.48 45.11
FiQA2018 16.99 24.62 27.24
HotpotQA 22.64 48.46 54.54
ImdbClassification 71.92 75.68 77.95
MSMARCO 07.03 18.81 19.13
MTOPDomainClassification 91.24 94.33 95.48
MTOPIntentClassification 74.08 79.54 82.84
MassiveIntentClassification 69.99 73.84 76.65
MassiveScenarioClassification 75.15 79.17 79.99
MedrxivClusteringP2P 30.15 30.94 30.11
MedrxivClusteringS2S 26.96 28.04 26.93
MindSmallReranking 29.52 30.86 29.73
NFCorpus 15.73 26.81 27.16
NQ 17.96 33.21 34.16
QuoraRetrieval 78.23 86.15 84.40
RedditClustering 38.67 42.84 41.83
RedditClusteringP2P 53.42 60.10 62.08
SCIDOCS 05.53 10.00 15.35
SICK-R 69.34 71.77 75.55
STS12 60.09 65.39 67.65
STS13 72.52 79.26 83.90
STS14 66.70 72.98 76.97
STS15 77.69 82.72 83.80
STS16 75.94 81.02 81.91
STS17 81.67 86.70 85.58
STS22 63.70 63.47 65.93
STSBenchmark 73.36 78.32 80.42
SciDocsRR 67.76 77.62 77.81
SciFact 38.31 64.48 68.67
SprintDuplicateQuestions 77.36 87.57 91.30
StackExchangeClustering 59.35 65.12 67.34
StackExchangeClusteringP2P 31.47 33.61 34.50
StackOverflowDupQuestions 40.82 47.77 49.80
SummEval 31.23 31.38 30.19
TRECCOVID 56.04 60.67 55.66
Touche2020 19.17 10.18 06.54
ToxicConversationsClassification 68.41 71.81 70.71
TweetSentimentExtractionClassification 56.08 57.17 60.90
TwentyNewsgroupsClustering 31.54 30.76 30.26
TwitterSemEval2015 61.54 65.14 68.76
TwitterURLCorpus 77.73 80.94 82.76
Average 49.42 55.36 56.80
Table 11: Unsupervised results of LLM2Vec transformed models on MTEB.
Task S-LLaMA-1.3B LLaMA-2-7B Mistral-7B
AmazonCounterfactualClassification 77.42 82.22 77.58
AmazonPolarityClassification 82.05 89.69 91.12
AmazonReviewsClassification 40.81 48.47 49.97
ArguAna 51.66 56.53 57.48
ArxivClusteringP2P 43.47 43.14 42.81
ArxivClusteringS2S 39.85 42.38 44.24
AskUbuntuDupQuestions 60.71 63.13 63.98
BIOSSES 85.88 82.13 85.24
Banking77Classification 86.01 88.17 88.31
BiorxivClusteringP2P 37.10 35.88 34.27
BiorxivClusteringS2S 34.28 34.81 35.53
CQADupstackRetrieval 41.73 45.94 48.84
ClimateFEVER 33.49 30.70 35.19
DBPedia 43.58 48.42 49.58
EmotionClassification 48.38 51.71 52.05
FEVER 86.81 89.93 89.40
FiQA2018 41.00 51.28 53.11
HotpotQA 63.85 72.99 74.07
ImdbClassification 75.33 85.78 87.42
MSMARCO 38.32 41.45 42.17
MTOPDomainClassification 94.09 95.57 96.04
MTOPIntentClassification 77.05 82.81 84.77
MassiveIntentClassification 75.58 78.06 79.29
MassiveScenarioClassification 79.16 81.35 81.64
MedrxivClusteringP2P 33.55 32.23 31.07
MedrxivClusteringS2S 31.11 31.37 31.27
MindSmallReranking 31.96 31.34 31.50
NFCorpus 37.12 40.33 39.33
NQ 53.89 61.24 61.70
QuoraRetrieval 87.37 85.59 87.75
RedditClustering 53.02 61.10 60.24
RedditClusteringP2P 60.47 64.52 64.12
SCIDOCS 17.96 21.05 22.50
SICK-R 82.25 83.01 83.70
STS12 78.28 78.85 78.80
STS13 85.52 86.84 86.37
STS14 82.49 84.04 84.04
STS15 88.76 88.72 88.99
STS16 87.11 86.79 87.22
STS17 90.10 90.63 90.19
STS22 68.25 67.55 67.68
STSBenchmark 87.16 88.72 88.65
SciDocsRR 79.23 84.03 83.80
SciFact 72.08 77.30 78.86
SprintDuplicateQuestions 96.25 96.83 96.82
StackExchangeClustering 63.04 67.98 70.73
StackExchangeClusteringP2P 34.01 33.20 34.50
StackOverflowDupQuestions 49.61 51.02 54.41
SummEval 30.01 28.49 29.96
TRECCOVID 80.41 79.25 77.69
Touche2020 22.31 16.92 22.18
ToxicConversationsClassification 69.92 71.01 69.26
TweetSentimentExtractionClassification 60.76 61.11 62.14
TwentyNewsgroupsClustering 49.37 51.04 52.18
TwitterSemEval2015 76.14 80.70 80.60
TwitterURLCorpus 86.23 86.56 86.56
Average 61.85 64.14 64.80
Table 12: Supervised results of LLM2Vec (only Bi + MNTP) models on MTEB.
We present the detailed performance of supervised LLM2Vec-transformed models on full MTEB in Table 12. Here, we only report the Bi + MNTP transformed models as we showed they perform the best after supervised fine-tuning.